Carland, Inc., has a project available with the following cash flows. If the required return for the project is 8.2 percent, what is the project's NPV?

Year Cash Flow
0 −$261,000
1 69,900
2 93,700
3 119,300
4 72,100
5 −12,300
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=69,900/1.082+93700/1.082^2+119300/1.082^3+72100/1.082^4

=291423.43

Present value of outflows=261000+12300/1.082^5

=269294.09

NPV=Present value of inflows-Present value of outflows

=291423.43-269294.09

=$22129.34(Approx).